# Memento-II: Learning by Stateful Reflective Memory

### v1

@@ 3.2.2 Transforming to Reflected MDP
To recover a Markovian formulation, we redefine the state to include memory, forming an augmented state $x_{t}=(s_{t},M_{t})$. The resulting process over $x_{t}$} is Markovian in this expanded space, even though the underlying environment process over $s_{t}$} is not. Each augmented state $x_{t}=(s_{t},M_{t})$ combines the environment state $s_{t}$ with the episodic memory $M_{t}$, as illustrated by the dashed rectangles in Figure 3.
In addition, since our focus is on the retrieval process rather than on modifying the LLM itself, we treat the LLM as a fixed component of the environment. Its action generation is absorbed into the environment’s transition and reward functions, effectively redefining the interaction dynamics. Under this view, the agent’s controllable decision is the retrieval action $c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$ while the environment, augmented by the fixed LLM kernel ($p_{\mathrm{LLM}}(a\mid s_{t},c_{t})$), determines both the resulting state transition and reward. This formulation leads naturally to a new MDP in which the retrieval policy ($\mu$) becomes the agent’s primary policy, and the LLM-mediated dynamics are incorporated into the environment model. Overall, we have our new MDP that is simplified by introducing an augmented state space and LLM becoming part of environment, namely,
@@
Given an underlying SRDP
$\mathcal{M}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle,$
with state space $\mathcal{S}$, action space $\mathcal{A}$, transition kernel $\mathcal{P}$, reward function $\mathcal{R}$, discount factor $\gamma$, episodic memory $\mathcal{M}$, and an LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$, the Reflected MDP is a tuple
$\mathcal{M}_{\mathrm{LLM}}=\langle\mathcal{X},\mathcal{C},\mathcal{P}_{\mathrm{LLM}},\mathcal{R}_{\mathrm{LLM}},\gamma\rangle,$
where the components are defined as follows:
  • Augmented state space: $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, with each state $x=(s,M)$ consisting of the current environment state $s$ and the current episodic memory $M$.
  • Action space: $\mathcal{C}=M$, corresponding to retrieval actions (selecting cases from memory).
@@
The Reflected MDP provides a fully Markovian formulation over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathcal{M}$, where transitions and rewards correspond to the expected outcomes under the LLM kernel $p_{\mathrm{LLM}}$. Each transition involves four steps: (1) retrieving a case $c_{t}\sim\mu(\cdot\mid s_{t},M_{t})$ from memory, (2) generating an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$, (3) sampling the next state $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$, and (4) updating the memory $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},\mathcal{R}(s_{t},a_{t}))$.
@@ 3.3 Policy Iteration
In order to enable an LLM agent to act even when no relevant past cases exist in memory, we augment the case set $\mathcal{C}$ with a special void case:
$\mathcal{M}_{v}\;=\;\mathcal{M}\cup\{c_{v}\}.$
The void case $c_{v}$ represents the ability of the agent to generate an action directly from its internal world knowledge, independently of retrieved memory. We formulate this by modifying the prior policy $\mu_{0}$ that incorporates a state-dependent mixture between memory-based retrieval and the void case. Let $K(x,c)$ be a similarity kernel for ordinary cases $c\in\mathcal{C}$, and let $K_{v}>0$ be a constant kernel score assigned to the void case $c_{v}$. Then, the prior distribution $\mu_{0}(\cdot\mid x)$ becomes
$\mu_{0}(c\mid x)\equiv\frac{\begin{cases}K(x,c),&c\neq c_{v},\\[3.0pt] K_{v},&c=c_{v},\end{cases}}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})+K_{v}}.$ (7)
@@
$\mu_{\mathrm{mem}}(c\mid x)=\frac{K(x,c)}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})},\qquad c\neq c_{v},$
@@
$\lambda(x)=\frac{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})}{\sum_{c^{\prime}\neq c_{v}}K(x,c^{\prime})+K_{v}}.$
@@
$\mu_{0}(c\mid x)=\lambda(x)\,\mu_{\mathrm{mem}}(c\mid x)+\bigl(1-\lambda(x)\bigr)\,\delta_{c_{v}}(c),$ (8)
where $\delta_{c_{v}}$ denotes the unit point mass at $c_{v}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{v}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [berry1997bandit], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
@@
$(\mathcal{T}^{\mu}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\left(\sum_{c^{\prime}}\mu(c^{\prime}\mid x^{\prime})\,Q(x^{\prime},c^{\prime})-\alpha\,\mathrm{KL}\!\bigl(\mu(\cdot\mid x^{\prime})\,\|\,\mu_{0}(\cdot\mid x^{\prime})\bigr)\right)\,\middle|\,x,c\right].$ (9)
@@
$\mu^{+}(c\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c\in M}\nu(c|x)\,Q(x,c)-\alpha\,\mathrm{KL}(\nu(c|x)\|\mu_{0}(\cdot\mid x))\right\}$ (10)
@@
$\mu^{+}(c\mid x)=\frac{\mu_{0}(c\mid x)\,\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}\mid x)\,\exp(Q(x,c^{\prime})/\alpha)}$ (11)
The proof can be found in Appendix A. Because $\mu_{0}(c_{v}\mid x)>0$ for all $x$, the improved policy $\mu^{\prime}$ always assigns nonzero probability to the void case, preserving the agent’s capacity for discovery. To summary, for fixed retrieval policy $\mu$, we define soft state-action values with the Parzen prior incorporated into the evaluation operator, and the soft $Q$-function for retrieval policy $\mu$ satisfies:
$Q^{\mu}(x,c)$ $=\mathbb{E}\left[r+\gamma V^{\mu}(x^{\prime})\mid x,c\right]$ (12)
$V^{\mu}(x)$ $=\sum_{c}\mu(c\mid x)\,Q^{\mu}(x,c)-\alpha\,\mathrm{KL}(\mu(\cdot\mid x)\|\mu_{0}(\cdot\mid x)),$ (13)
where the expectation integrates over $a\sim p_{\mathrm{LLM}}(\cdot\mid s,c)$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s,a)$, and $M^{\prime}=\mathsf{Write}(M,s,a,Q)$.
KL-Regularised Policy Iteration Now, on the basis of policy evaluation in Eq. (9) and policy improvement in Eq. (11), we are ready to give the policy iteration algorithm, as shown in Algorithm 1.
Algorithm 1 Stateful Reflective Learning with Read–Write Loop
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$, initial $Q_{0}$, initial reference prior $\mu_{0,0}$ (with void case $c_{v}$), initial memory $M_{0}$
2:for $k=0,1,2,\ldots$ do
3:  (Prior Update) Using memory $M_{k}$, construct the Parzen kernel-based reference prior $\mu_{0,k+1}(\cdot\mid x)$ according to (6), including the void case $c_{v}$ via the mixture form in (8).
4:  (Memory Read, a.k.a, Policy Improvement) Update the policy by the KL-regularised greedy step
$\mu_{k+1}(\cdot\mid x)\propto\mu_{0,k+1}(\cdot\mid x)\,\exp\bigl(Q^{\mu_{k}}(x,\cdot)/\alpha\bigr),$
i.e. using the closed-form solution in (11).
5:  (Interact with Environment) Execute the policy $(\mu_{k+1},p_{\mathrm{LLM}})$ and receive feedback $r$.
6:  (Memory Rewrite, a.k.a., Policy Evaluation) Collect transitions $(s,a,r)$, and update memory
$M_{k+1}\leftarrow\mathsf{Write}(M_{k},s,a,r).$
7:   Given $\mu_{k}$ and $\mu_{0,k}$, approximately solve $Q^{\mu_{k}}$ by iterating
$Q\leftarrow\mathcal{T}^{\mu_{k}}_{\mathrm{KL}}Q$
using the KL-regularised evaluation operator in (9). Two classes of memory representations: (i) a tabular memory that explicitly stores $Q$-values for each state–action tuple, and (ii) a parameterised memory in which the $Q$-function is represented by a neural network.
@@
Under standard assumptions, this procedure yields a monotone improvement in the KL-regularised value and converges to the fixed point of the KL-regularised optimality operator,
Specifically, we define
@@
Under assumptions 5 and 6 (bounded rewards and $\gamma<1$ and stationary memory), the process (the reflected MDP) is an ordinary discounted MDP. Algorithm 1 converges to a fixed point $(Q^{\star},\mu^{\star})$ that is optimal for the KL-regularised objective with Parzen prior.
@@
Discussion It is worth mentioning that the void case $c_{v}$ plays a crucial role: by providing a state-independent baseline probability in $\mu_{0}$, it ensures that policy iteration retains the ability to discover new behaviours beyond exploitation or exploration of existing memory. Unlike classical density estimation, our memory cases are not independent and identically distributed but are gathered sequentially through policy execution. Moreover, the finite memory constraint means we perform kernel smoothing over a limited and evolving set of points rather than relying on asymptotic density estimation. As a result, bandwidth selection plays a critical role in balancing local similarity against statistical robustness as the memory continues to grow. Several strategies can be used to choose the bandwidth $h$. Silverman’s rule suggests setting $h\propto n^{-1/(d+4)}$, where $n=|M|$ is the memory size and $d$ the embedding dimension. Another approach is cross-validation, which tunes $h$ by optimising performance on held-out retrieval tasks. Finally, an adaptive strategy decreases $h$ as memory grows, ensuring that neighbourhoods remain local and informative.
3.4 Two-Time-Scale Convergence with Online Memory Re-writing
We next allow memory to evolve on a slower time-scale. Let $\eta_{t}$ be steps for evaluation/improvement and $\rho_{t}$ for memory updates, with $\rho_{t}/\eta_{t}\to 0$, $\sum_{t}\eta_{t}=\infty$, $\sum_{t}\eta_{t}^{2}<\infty$. We make the following assumptions.

### v2

@@ 3.2.2 Transforming to Reflected MDP
To recover a Markovian formulation, we redefine the state to explicitly include memory, forming an augmented state, i.e., at time $t$, the agent occupies an augmented state
$x_{t}:=(s_{t},M_{t})\in\mathcal{X}:=S\times\mathfrak{M}.$
The resulting process over $\{x_{t}\}$ is Markovian in this expanded space, even though the environment process over $\{s_{t}\}$ alone is not. Each augmented state combines the current environment state $s_{t}$ with the agent’s episodic memory $M_{t}$, as illustrated by the dashed rectangles in Figure 3. Augmenting the state with memory restores the Markov property and enables a well-defined control formulation, i.e., for any reflection action $c_{t}$,
$\Pr(x_{t+1}\mid x_{0:t},c_{0:t})=\Pr(x_{t+1}\mid x_{t},c_{t}),$
where the environment transition depends only on $(s_{t},a_{t})$, while the memory update is a deterministic or stochastic function of $(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$. Since the reflection mechanism and the LLM action distribution are fixed functions of $(s_{t},M_{t})$, all sources of dependence on the past are captured by $x_{t}$. Hence the augmented process satisfies the Markov property.
Since our focus is on learning through memory retrieval rather than modifying the LLM itself, we treat the LLM as a fixed or slow-evolving component of the environment. Its action generation is absorbed into the environment dynamics via the effective transition and reward functions $\mathcal{P}^{\mathrm{LLM}}$ and $\mathcal{R}^{\mathrm{LLM}}$. Under this view, the agent’s only controllable decision is the retrieval action
$c_{t}\sim\mu(\cdot\mid s_{t},M_{t}),$
while state transitions and rewards are determined jointly by the environment and the fixed LLM kernel $p_{\mathrm{LLM}}(a\mid s_{t},c_{t})$. The resulting Reflected MDP therefore treats the retrieval policy $\mu$ as the agent’s primary policy, with LLM mediated reasoning incorporated into the environment model. Specifically, we have:
@@
the Reflected MDP is a tuple
$\mathcal{D}_{\mathrm{ReMDP}}=\langle\mathcal{X},\mathcal{C},\mathcal{P}_{\mathrm{LLM}},\mathcal{R}_{\mathrm{LLM}},\gamma\rangle,$
that transforms the underlying SRDP, $\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathfrak{M},p_{\mathrm{LLM}}\rangle$, with the components are defined as follows:
  • Augmented state space: $\mathcal{X}=\mathcal{S}\times\mathfrak{M}$, with each state $x=(s,M)$ consisting of the current environment state $s$ and the current episodic memory $M$.
  • Action space: $\mathcal{C}(M):=M$, consisting of retrieval actions that index and select cases from the current episodic memory.
@@
The Reflected MDP provides a fully Markovian formulation over the augmented state space $\mathcal{X}=\mathcal{S}\times\mathfrak{M}$, in which transitions and rewards correspond to expected outcomes under the fixed LLM kernel $p_{\mathrm{LLM}}$. At each time step, the transition unfolds as follows: the agent first retrieves a memory case $c_{t}\sim\mu(\cdot\mid s_{t},M_{t})$; conditioned on this retrieval, the LLM generates an environment action $a_{t}\sim p_{\mathrm{LLM}}(\cdot\mid s_{t},c_{t})$; the environment then transitions to $s_{t+1}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$ and produces a reward $\mathcal{R}(s_{t},a_{t})$; finally, episodic memory is updated according to $M_{t+1}=\mathsf{Write}(M_{t},s_{t},a_{t},r_{t},s^{\prime}_{t})$.
By absorbing the retrieval and LLM generation mechanisms into effective transition and reward functions over the augmented state space $\mathcal{X}$, this interaction induces a simplified control process. A policy in the Reflected MDP is defined as a retrieval policy
$\mu:\mathcal{X}\to\Delta(\mathcal{C}),$
which maps augmented states x=(s,M)x=(s,{M}) to probability distributions over retrieval actions. Environment actions are generated implicitly by the fixed LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$ and absorbed into the transition dynamics. This abstraction captures the full reflective behaviour of the agent in a compact Markovian form that is amenable to classical reinforcement learning analysis and learning algorithms.
It is worth mentioning that this augmented state admits a natural cognitive interpretation. The environment state $s_{t}$ encodes the agent’s external sensory or task-level signal, while the episodic memory $M_{t}$ represents an internal cognitive state that stores and organises past experiences, outcomes, and contextual information. The pair $(s_{t},M_{t})$ therefore constitutes the agent’s effective decision state, integrating external observation with internal experience. This perspective parallels belief-state constructions in partially observable Markov decision processes, where internal memory summarises relevant history to support optimal control [kaelbling1998planning], and aligns with cognitive theories that view episodic memory as a core substrate for human decision making and learning [tulving1983elements, GershmanDaw2017EpisodicRL].
The underlying MDP captures the dynamics of the environment alone, independent of the agent’s internal mechanisms. In contrast, the SRDP explicitly models the coupled interaction among the agent, its episodic memory, and the LLM during decision making. The Reflected MDP then provides a Markovian reformulation of this process by lifting the SRDP to an augmented state space, in which the combined agent–memory–LLM dynamics become Markovian. The relationship between the underlying environment, the SRDP, and the resulting Reflected MDP is summarised as
$\underbrace{\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\rangle}_{\text{Underlying MDP}}\xrightarrow{\;\mathrm{ext.}\;}\underbrace{\Big\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,M,\mu,p_{\mathrm{LLM}}\Big\rangle}_{\text{SRDP}}\xrightarrow{\;\mathrm{equiv.}\;}\underbrace{\Big\langle\mathcal{X}:=\mathcal{S}\times M,\mathcal{C},\mathcal{P}^{\mathrm{LLM}},\mathcal{R}^{\mathrm{LLM}},\gamma\Big\rangle}_{\text{Reflected MDP}},$ (4)
where a detailed comparison between the SRDP and the Reflected MDP is provided in Appendix E.
4 Read–Write Reflective Learning
Under the reflected MDP abstraction, continual learning reduces to an iterative read–write reflective procedure over episodic memory, which we formalise as a learning algorithm next.
@@ 4.1 Policy Iteration
In order to enable an LLM agent to act even when no relevant past cases exist in memory, we augment the case set $\mathcal{C}$ with a special void case: $M\leftarrow M\cup\{c_{\varnothing}\}$. The void case $c_{\varnothing}$ represents the ability of the agent to generate an action directly from its internal world knowledge, independently of retrieved memory. We formulate this by modifying the prior policy $\mu_{0}$ that incorporates a state-dependent mixture between memory-based retrieval and the void case. Let $K(x,c)$ be a similarity kernel for ordinary cases $c\in\mathcal{C}$, and let $K_{\varnothing}>0$ be a constant kernel score assigned to the void case $c_{\varnothing}$. Then, the prior distribution $\mu_{0}(\cdot\mid x)$ becomes
$\mu_{0}(c\mid x)\equiv\frac{\begin{cases}K(x,c),&c\neq c_{\varnothing},\\[3.0pt] K_{\varnothing},&c=c_{\varnothing},\end{cases}}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})+K_{\varnothing}},$ (7)
where $K(x,c):=K\!\left(\frac{\psi(s(x))-\psi(s(c))}{h}\right)$.
@@
$\mu_{\mathrm{mem}}(c\mid x)=\frac{K(x,c)}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})},\qquad c\neq c_{\varnothing},$
@@
$\lambda(x)=\frac{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})}{\sum_{c^{\prime}\neq c_{\varnothing}}K(x,c^{\prime})+K_{\varnothing}}.$
@@
$\mu_{0}(c\mid x)=\lambda(x)\,\mu_{\mathrm{mem}}(c\mid x)+\bigl(1-\lambda(x)\bigr)\,\delta_{c_{\varnothing}}(c),$ (8)
where $\delta_{c_{\varnothing}}$ denotes the unit point mass at $c_{\varnothing}$. This mixture form shows that the agent automatically interpolates between reflection(retrieval)-driven behaviour when similarity to stored cases is high ($\lambda(x)\approx 1$) and knowledge-driven discovery through $c_{\varnothing}$ when similarity is low ($\lambda(x)\approx 0$). This mechanism mirrors the “discovery” component in infinite-armed bandits [berry1997bandit], complementing exploitation and exploration with the capacity to act beyond the current memory. This mechanism ensures that the agent can always act, even in novel states, while still exploiting relevant past experiences when available.
@@
$(\mathcal{T}^{\mu}_{\mathrm{KL}}Q)(x,c):=\mathbb{E}_{a,s^{\prime},M^{\prime}}\!\left[r+\gamma\left(\sum_{c^{\prime}}\mu(c^{\prime}\mid x^{\prime})\,Q(x^{\prime},c^{\prime})-\alpha\,\mathrm{KL}\!\bigl(\mu(\cdot\mid x^{\prime})\,\|\,\mu_{0}(\cdot\mid x^{\prime})\bigr)\right)\,\middle|\,x,c\right],$ (9)
where $\mathcal{T}^{\mu}_{\mathrm{KL}}:\mathbb{R}^{\mathcal{X}\times\mathcal{C}}\to\mathbb{R}^{\mathcal{X}\times\mathcal{C}}$. It is worth emphasising that all value functions $Q(x,c)$ and $V(x)$ defined on the augmented state space $\mathcal{X}$ of the Reflected MDP are distinct from the value functions $Q^{\mathcal{D}}$ and $V^{\mathcal{D}}$ of the original MDP (see Section 3.1). The latter are defined over the environment state space $S$, whereas the former explicitly incorporate episodic memory and LLM-mediated action generation through the Reflected MDP dynamics. We shall show that, by appropriately leveraging these LLM-induced dynamics, the Reflected MDP can approximate the optimal value functions of the original MDP in a more sample-efficient manner.
@@
$\mu^{+}(c\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c\in M}\nu(c|x)\,Q(x,c)-\alpha\,\mathrm{KL}(\nu(c|x)\|\mu_{0}(\cdot\mid x))\right\},$ (10)
@@
$\mu^{+}(c\mid x)=\frac{\mu_{0}(c\mid x)\,\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in M}\mu_{0}(c^{\prime}\mid x)\,\exp(Q(x,c^{\prime})/\alpha)}.$ (11)
The proof can be found in Appendix A. Because $\mu_{0}(c_{\varnothing}\mid x)>0$ for all $x$, the improved policy $\mu^{\prime}$ always assigns nonzero probability to the void case, preserving the agent’s capacity for discovery.
KL-Regularised Policy Iteration Now, on the basis of policy evaluation in Eq. (9) and policy improvement in Eq. (11), we are ready to give the policy iteration algorithm, as shown in Algorithm 1. Under standard assumptions, this procedure yields a monotone improvement in the KL-regularised value and converges to the fixed point of the KL-regularised optimality operator,
Algorithm 1 Soft Policy Iteration with Fixed Memory under Reflected MDP
1:Input: Temperature $\alpha>0$, kernel bandwidth $h>0$
2:Initialise: Fixed episodic memory $M$, reference prior $\mu_{0}(\cdot\mid x)$ (including void case $c_{\varnothing}$), initial policy $\mu_{0}$, initial $Q_{0}$
3:for $t=0,1,2,\ldots$ do
4:  (Policy Evaluation)
5:  Compute the KL-regularised value function $Q^{\mu_{t}}$ as the unique fixed point of
$Q=\mathcal{T}^{\mu_{t}}_{\mathrm{KL}}Q,$
where (Eq. (9))
$(\mathcal{T}^{\mu_{t}}_{\mathrm{KL}}Q)(x,c)=\mathbb{E}\!\left[r+\gamma\Big(\sum_{c^{\prime}\in M}\mu_{t}(c^{\prime}\mid x^{\prime})Q(x^{\prime},c^{\prime})-\alpha\mathrm{KL}(\mu_{t}(\cdot\mid x^{\prime})\|\mu_{0}(\cdot\mid x^{\prime}))\Big)\,\middle|\,x,c\right].$
6:  (Policy Improvement)
7:  Update the retrieval policy by solving
$\mu_{k+1}(\cdot\mid x)\in\arg\max_{\nu\in\Delta(M)}\left\{\sum_{c}\nu(c\mid x)Q^{\mu_{t}}(x,c)-\alpha\mathrm{KL}(\nu(\cdot\mid x)\|\mu_{0}(\cdot\mid x))\right\},$
which admits the closed-form solution (Eq. (11))
$\mu_{t+1}(c\mid x)=\frac{\mu_{0}(c\mid x)\exp(Q^{\mu_{t}}(x,c)/\alpha)}{\sum_{c^{\prime}}\mu_{0}(c^{\prime}\mid x)\exp(Q^{\mu_{t}}(x,c^{\prime})/\alpha)}.$
@@
For the convergence analysis, we start with assumptions followed with the theorem:
@@
Under assumptions 6 and 7 (bounded rewards and $\gamma<1$ and stationary memory), the process (the reflected MDP) is an ordinary discounted MDP. Iterations between (9) and (11) (Algorithm 1) converge to a fixed point $(Q^{\star},\mu^{\star})$ that is optimal for the KL-regularised objective with Parzen prior.
@@
Some discussions: It is worth mentioning that the void case $c_{\varnothing}$ plays a crucial role: by providing a state-independent baseline probability in $\mu_{0}$, it ensures that policy iteration retains the ability to discover new behaviours beyond exploitation or exploration of existing memory. Unlike classical density estimation, our memory cases are not independent and identically distributed but are gathered sequentially through policy execution. Moreover, the finite memory constraint means we perform kernel smoothing over a limited and evolving set of points rather than relying on asymptotic density estimation. As a result, bandwidth selection plays a critical role in balancing local similarity against statistical robustness as the memory continues to grow. Several strategies can be used to choose the bandwidth $h$. Silverman’s rule suggests setting $h\propto n^{-1/(d+4)}$, where $n=|M|$ is the memory size and $d$ the embedding dimension. Another approach is cross-validation, which tunes $h$ by optimising performance on held-out retrieval tasks. Finally, an adaptive strategy decreases $h$ as memory grows, ensuring that neighbourhoods remain local and informative.
4.2 Two-Time-Scale Convergence with Online Memory Rewriting
The preceding analysis assumes that episodic memory remains fixed. We now relax this assumption and allow memory to evolve on a slower time scale. Let $\eta_{t}$ denote the step size used for policy evaluation and improvement, and let $\rho_{t}$ denote the step size for memory updates. We assume a two time scale regime in which
$\frac{\rho_{t}}{\eta_{t}}\to 0,\qquad\sum_{t=1}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty.$
Under this separation of time scales, policy evaluation and improvement operate on a faster time scale, while episodic memory evolves more slowly. Formally, we make the following assumptions: